OpenStudy (kayla254):

jake is comparing the prices of two mattress cleaning companies. company A charges 30 per matress and an additional 13 as sevice changes. company B charges 28 per matress and an additional 15 as service changes. part A : write equations to represent company a's and company b's total matress cleaning charges for a certian number of matresses .define the variable used in the equations .Part B:which company would charge less for cleaning 4 matresses.Part C:how much money is saved by using the services of company B instead on company A to clean 7 matresses

OpenStudy (anonymous):

Hello, welcome to OpenStudy. This is the OpenStudy Feedback section, and I think that your question would be better suited for the Mathematics section. openstudy.com/study#/groups/Mathematics for reference next time. I help you out anyway though as you are new here.

OpenStudy (anonymous):

So, Part A requires to different equations, one for company A, and one for company B. Company A: $30 per mattress, + 13 charge So, we need a variable to represent the total cost, and a variable to represent the number of mattresses. We'll use C for cost, and m for mattress. C=30m+13 Company B will be very similar, with minor adjustments to the fixed costs. C=28m+15

OpenStudy (anonymous):

Part B requires us to figure out which company will charge less for four mattresses. I can already tell you it will be company B because of the reference in Part C, but let's explain why with mathematical reasoning. We'll plug 4 in for m because m represents the number of mattresses. C=30(4)+13 C=120+13 C=133 C=28(4)+15 C=112+15 C=127 So, company B is cheaper when cleaning 4 mattresses.

OpenStudy (anonymous):

Part C requires us to find the difference in price between the companies if they were to each clean 7 mattresses. C=30(7)+13 C=210+13 C=223 ------------ C=28(7)+15 C=196+15 C=211 It wants to know how much is saved, so 223-211=12 So, $12 was saved by using company B.

OpenStudy (anonymous):

Feel free to tag me @OrangeMaster if you don't understand a part of the solution. @kayla254

OpenStudy (kayla254):

OMG THANKYOU SO MUCH @OrangeMaster

OpenStudy (anonymous):

My pleasure. :)

OpenStudy (kayla254):

@OrangeMaster i have one more question please help , PartA; solve -np-80<60 for n . show work please PartB: solve 2a -5d=30 for D.show how you got the answer

OpenStudy (anonymous):

If you'd like to close this question, and open your two new questions in the mathematics section. I'd be glad to help :)

OpenStudy (kayla254):

ok